# Mobility Edge for the Anderson Model on Random Regular Graphs

### 15 Mar 2026 v1

@@ 4.4. Conclusion
Fix $s\in(0,1/2)$ and $\delta\in(0,\eta^{-1}]$. Then for every $u>0$,
$\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\big]\big|\geq u\Big)\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Cdt^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big).$ (4.11)
@@
Let $s\in(0,1/2)$ and $\delta\in(0,\eta^{-1}]$. Then for every $u>0$,
@@
Next, by Lemma 4.10 and Lemma 4.9, for any $s\in(0,1/2)$, $\delta\in(0,\eta^{-1}]$, and $u>0$,
@@
$\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Cdt^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big),$
@@
5. Bounds for the Eigenvalue Counting Function
This section is dedicated to the proof of Lemma 3.4. To obtain this result, we first show that the Stieltjes transform of $W_{N}$ converges to the Stieltjes transform of the density of the local states measure of $H_{d,t}$ at the root of the Bethe lattice. For the Anderson model on the random regular graph with a bounded potential, an analogous result for the integrated density of states was proved by Geisinger [23]. We adapt this result to $W_{N}$ by a truncation argument, then upgrade the convergence in means to convergence in probability using the $s=1$ case of Lemma 4.1. Finally, since convergence of Stieltjes transforms implies convergence of the associated spectral measures, the desired lower bound on the eigenvalue counting bound follows from a lower bound on the integrated density of states of $H_{d,t}$, which we show in Appendix B, and the upper bound follows from Wegner’s estimate.
5.1. Notation
For all $x\in[N]$, we define the spectral measure of $W_{N}$ at $x$ by
$\mu_{N,x}(B)=\langle\delta_{x},\mathbbm{1}_{B}(W_{N})\delta_{x}\rangle,$
where $B\subset\mathbb{R}$ is any Borel set, and $\mathbbm{1}_{B}(W_{N})$ is defined in the usual way (using the spectral theorem). We define
$m_{N}(z)=\frac{1}{N}\operatorname{Tr}G_{N}(z),$
Next, bcause $H_{d,t}$ is self-adjoint, the spectral theorem implies the existence of a random projection-valued measure $\mu$ such that $H=\int\lambda\,d\mu(\lambda)$. The spectral measure at the root is given by
$\mu_{0}(S)=\langle\delta_{0},\mu(S)\delta_{0}\rangle$
$\nu(S)=\mathbb{E}\big[\mu_{0}(S)\big].$
For all $z\in\mathbb{H}$, the Stieltjes transform of $\nu$ is
$m_{\mathbb{T}}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,d\nu(\lambda).$
5.2. Preliminaries
We first record an integration by parts bound for signed distribution functions.
Lemma 5.1.
Let $\mu_{1}$ and $\mu_{2}$ be finite Borel measures on $\mathbb{R}$, and set
$F(t)=\mu_{1}\big((-\infty,t]\big)-\mu_{2}\big((-\infty,t]\big).$
If $f\in C^{1}(\mathbb{R})$ satisfies $f(\lambda)\to 0$ as $|\lambda|\to\infty$ and $f^{\prime}\in L^{1}(\mathbb{R})$, then
$\Big|\int_{\mathbb{R}}f\,d\mu_{1}-\int_{\mathbb{R}}f\,d\mu_{2}\Big|\leq\|F\|_{\infty}\int_{\mathbb{R}}\big|f^{\prime}(\lambda)\big|\,d\lambda.$
Proof.
We apply integration by parts to the signed measure $\mu_{1}-\mu_{2}$ (see, e.g., [21, Theorem 3.36]), yielding
$\int_{\mathbb{R}}f\,d(\mu_{1}-\mu_{2})=-\int_{\mathbb{R}}F(\lambda)f^{\prime}(\lambda)\,d\lambda,$
then bound the right-hand side by by $\|F\|_{\infty}\int|f^{\prime}|$. ∎
Next we establish an estimate on the stability of resolvents under perturbations.
Lemma 5.2.
Let $H$ and $\widetilde{H}$ be symmetric $N\times N$ matrices, and set $\Delta=\widetilde{H}-H$. Then for any $z\in\mathbb{H}$,
$\Big|\frac{1}{N}\operatorname{Tr}(\widetilde{H}-z)^{-1}-\frac{1}{N}\operatorname{Tr}(H-z)^{-1}\Big|\leq\frac{1}{N\eta^{2}}\operatorname{Tr}|\Delta|.$ (5.1)
Proof.
By the definition of a resolvent,
$(\widetilde{H}-z)^{-1}-(H-z)^{-1}=-(H-z)^{-1}\Delta(\widetilde{H}-z)^{-1}.$
Taking traces and using the trace norm bound $|\operatorname{Tr}(ABC)|\leq\|A\|\|C\|\operatorname{Tr}|B|$ yields
$\Big|\operatorname{Tr}\big((\widetilde{H}-z)^{-1}-(H-z)^{-1}\big)\Big|\leq\big\|(H-z)^{-1}\big\|\big\|(\widetilde{H}-z)^{-1}\big\|\operatorname{Tr}|\Delta|\leq\eta^{-2}\operatorname{Tr}|\Delta|.$
Dividing by $N$ yields the claim. ∎
Given a probability density $\tilde{\rho}$, we let $\widetilde{W}_{N}=-tA_{N}+\widetilde{V}_{N}$ and $H_{d,t}=-tA^{(d)}+\widetilde{V}^{(d)}$, where $\widetilde{V}_{N}$ and $\widetilde{V}^{(d)}$ are defined analogously to $V_{N}$ and $V^{(d)}$ with the Gaussian potential variables replaced by i.i.d. variables with density $\tilde{\rho}$. We set $\tilde{\mu}_{N,x}$ and $\tilde{\mu}_{0}$ denote the corresponding spectral measures, defined analogously to $\mu_{N,x}$ and $\mu_{0}$.
The following theorem is [23, Theorem 2.3].
Theorem 5.3.
Suppose that $\|\tilde{\rho}\|$$\infty$ and there exists $\rho_{0}>0$ such that $\mathrm{supp}(\tilde{\rho})\subset(-\rho_{0},\rho_{0})$. Let $\widetilde{\mathbb{E}}$ denote expectation with respect to the potentials $\tilde{V}_{N}$ and $\tilde{V}^{(d)}$, conditional on $A_{N}$. Define the event
$\mathcal{A}_{N}=\left\{\sup_{s\in\mathbb{R}}\Bigg|\frac{1}{N}\sum_{x\in[N]}\widetilde{\mathbb{E}}\big[\tilde{\mu}_{N,x}\big((-\infty,s]\big)\big]-\widetilde{\mathbb{E}}\big[\tilde{\mu}_{0}\big((-\infty,s]\big)\big]\Bigg|\leq\frac{5\pi\|\tilde{\rho}\|_{\infty}(2\sqrt{d-1}+\rho_{0})}{\operatorname{log}_{d-1}N}\right\},$ (5.2)
which depends only on the randomness in $A_{N}$. Then $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{A}_{N})=1$.
Define the disorder-averaged measures
$\tilde{\nu}^{*}_{N}=\frac{1}{N}\sum_{x\in[N]}\widetilde{\mathbb{E}}[\tilde{\mu}_{N,x}],\qquad\tilde{\nu}^{*}=\widetilde{\mathbb{E}}[\tilde{\mu}_{0}].$
The corresponding Stieltjes measures are
$m^{*}_{N}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,\tilde{\nu}^{*}_{N}(d\lambda),\qquad m^{*}_{\mathbb{T}}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,\tilde{\nu}^{*}(d\lambda).$
Using Theorem 5.3, we deduce convergence of the disorder-averaged Stieltjes transform in the bounded potential setting.
Proposition 5.4.
Retain the hypotheses and notation of Theorem 5.3. For every $z\in\mathbb{H}$, define the event
$\mathcal{B}_{N}(z)=\left\{\big|m^{*}_{N}(z)-m^{*}_{\mathbb{T}}(z)\big|\leq\frac{\pi}{\eta}\cdot\frac{5\pi\|\tilde{\rho}\|_{\infty}(2\sqrt{d-1}+\rho_{0})}{\operatorname{log}_{d-1}N}\right\},$ (5.3)
which depends only on the randomness in $A_{N}$. Then $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{B}_{N}(z))=1$.
Proof.
Define $f_{z}(\lambda)=(\lambda-z)^{-1}$. Then $f^{\prime}_{z}(\lambda)=-(\lambda-z)^{-2}$ and
$\int_{\mathbb{R}}|f^{\prime}_{z}(\lambda)|\,d\lambda=\int_{\mathbb{R}}\frac{d\lambda}{(\lambda-E)^{2}+\eta^{2}}=\frac{\pi}{\eta}.$ (5.4)
Apply Lemma 5.1 with $\mu_{1}=\tilde{\nu}^{*}_{N}$, $\mu_{2}=\tilde{\nu}^{*}$, and $f=f_{z}$. Further, define
$F_{N}(s)=\tilde{\nu}^{*}_{N}\big((-\infty,s]\big)\qquad F_{\mathbb{T}}(s)=\tilde{\nu}^{*}\big((-\infty,s]\big).$
Then, using (5.4),
$\big|m^{*}_{N}(z)-m^{*}_{\mathbb{T}}(z)\big|=\Big|\int f_{z}\,d\tilde{\nu}^{*}_{N}-\int f_{z}\,d\tilde{\nu}^{*}\Big|\leq\|F_{N}-F_{\mathbb{T}}\|_{\infty}\int|f^{\prime}_{z}|=\|F_{N}-F_{\mathbb{T}}\|_{\infty}\cdot\frac{\pi}{\eta}.$
Now insert the bound (5.2) to complete the proof. ∎
Next, we state an upper bound on the integrated density of states of $H_{d,t}$, which is a consequence of a well-known estimate of Wegner [8, (5.27)].
Lemma 5.5.
Fix $g,\mathfrak{L}>0$ such that (2.1) holds. For every $d\geq 3$, there exists a constant $C(d,\mathfrak{L})>0$ such that for all intervals $I\subset\mathbb{R}$, we have $\nu(I)\leq C|I|$.
Finally, we state a lower bound that complements the previous Wegner estimate. It is proved in Appendix B.
Lemma 5.6.
Fix $g,\mathfrak{L}>0$ such that (2.1) holds, and fix $D>0$. For every $d\geq 3$, there exists a constant $c(d,D,\mathfrak{L})>0$ such that for all intervals $I\subset[-D,D]$, we have $\nu(I)\geq c|I|$.
5.3. Global Law for Gaussian Disorder
To pass from Gaussian disorder to bounded support, we introduce a coupled truncation scheme. Fix $L\geq 1$. Let $X\sim N(0,1)$ and let $U^{(L)}\sim\mathrm{Unif}[-L,L]$ be independent of $V$. Define
$X^{(L)}=X\mathbbm{1}_{\{|X|\leq L\}}+U^{(L)}\mathbbm{1}_{\{|X|>L\}}.$ (5.5)
Let $(X_{x},U_{x}^{(L)})_{x\in[N]}$ be i.i.d. with law $(X,U)$, define $X_{x}^{(L)}$ by (5.5), and set
$V_{N}^{(L)}=\mathrm{diag}(X_{1}^{(L)},\dots,X_{N}^{(L)}),\qquad W_{N}^{(L)}=-tA_{N}+V_{N}^{(L)}.$
Define the truncated Stieltjes transform
$m_{N}^{(L)}(z)=\frac{1}{N}\operatorname{Tr}(W_{N}^{(L)}-z)^{-1},$
and define similarly $H_{t,d}^{(L)}=-tA^{(d)}+V^{(d),(L)}$ on the Bethe lattice and
$m_{\mathbb{T}}^{(L)}(z)=\mathbb{E}\big[\langle\delta_{0},(H_{d,t}^{(L)}-z)^{-1}\delta_{0}\rangle\big].$
Proposition 5.7.
Proof.
We proceed through a sequence of reductions that isolate truncation and approximation errors. We first apply Lemma 5.2 with $H=W_{N}$ and $\widetilde{H}=W_{N}^{(L)}$. Here $\Delta$ is diagonal with entries $V_{N,i}^{(L)}-V_{N,i}$, hence
$|m_{N}(z)-m_{N}^{(L)}(z)|\leq\frac{1}{\eta^{2}}\cdot\frac{1}{N}\sum_{i=1}^{N}|V_{N,i}^{(L)}-V_{N,i}|.$
Taking expectation and using that the entries of each potential are i.i.d., we find
$\sup_{N\geq 1}\mathbb{E}\Big[\big|m_{N}(z)-m_{N}^{(L)}(z)\big|\Big]\leq\frac{1}{\eta^{2}}\mathbb{E}\big[|V_{N,1}^{(L)}-V_{N,1}|\big]\leq\frac{1}{\eta^{2}}\mathbb{E}\big[\big(|g_{1}|+L\big)\mathbbm{1}_{\{|g_{1}|>L\}}\big].$ (5.6)
For each $L$ and $i\in[N]$, the law of $V^{(L)}_{N,i}$ is supported on [L,L][-L,L] and has a bounded density. Therefore Proposition 5.4 applies to $W_{N}^{(L)}$ and yields, for fixed $L$ and fixed $z$,
$\lim_{N\rightarrow\infty}\mathbb{E}\big[m_{N}^{(L)}(z)\big]=m_{\mathbb{T}}^{(L)}(z).$ (5.7)
We now remove the truncation on the tree. Because the tree has infinitely many sites, we do not compare $m_{\mathbb{T}}^{(L)}$ and $m_{\mathbb{T}}$ by summing over sites of $\mathbb{T}_{d}$. Instead, we use strong resolvent convergence. Let $C_{c}(\mathbb{T})$ denote finitely supported functions on $T_{d}$. It is well known that $C_{c}(\mathbb{T})$ is a core for the multiplication operator $V^{(d)}$ and hence for $H_{d,t}=-tA^{(d)}+V^{(d)}$ (since $tA^{(d)}$ is bounded on $\ell^{2}(\mathbb{T})$). Moreover, for each $\psi\in C_{c}(\mathbb{T})$, we have $H_{d,t}^{(L)}\psi\to H_{d,t}\psi$ a.s. as $L\to\infty$ because $\psi$ has finite support and $V_{x}^{(L)}\to V_{x}$ pointwise. By the core convergence criterion for strong resolvent convergence (see, e.g., [31, Theorem VIII.25]), $(H_{d,t}^{(L)}-z)^{-1}$ converges in the strong operator topology to $(H_{d,t}-z)^{-1}$. Applying this fact to $\delta_{0}$ and using the uniform bound $\|(H_{d,t}^{(L)}-z)^{-1}\|\leq\eta^{-1}$ together with the dominated convergence theorem yields
$\lim_{L\rightarrow\infty}m_{\mathbb{T}}^{(L)}(z)=m_{\mathbb{T}}(z).$ (5.8)
Finally, we combine the previous estimates to conclude convergence in mean, then upgrade to convergence in probability. Fix $\varepsilon>0$. Choose $L$ large enough so that (5.6) and (5.8) yield
$\sup_{n}\mathbb{E}\big[\big|m_{N}(z)-m_{N}^{(L)}(z)\big|\big]<\varepsilon/3,\qquad|m_{\mathbb{T}}^{(L)}(z)-m_{\mathbb{T}}(z)|<\varepsilon/3.$
Then take $N$ large enough so that (5.7) gives $|\mathbb{E}m_{N}^{(L)}(z)-m_{\mathbb{T}}^{(L)}(z)|<\varepsilon/3$. By the triangle inequality,
$\big|\mathbb{E}\big[m_{N}(z)\big]-m_{\mathbb{T}}(z)\big|\leq\big|\mathbb{E}\big[m_{N}(z)\big]-\mathbb{E}\big[m_{N}^{(L)}(z)\big]\big|+\big|\mathbb{E}\big[m_{N}^{(L)}(z)\big]-m_{\mathbb{T}}^{(L)}(z)\big|+\big|m_{\mathbb{T}}^{(L)}(z)-m_{\mathbb{T}}(z)\big|<\varepsilon.$
Hence
$\lim_{N\rightarrow\infty}\mathbb{E}\big[m_{N}(z)\big]=m_{\mathbb{T}}(z)$ (5.9)
in probability. By the $s=1$ case of Lemma 3.3, there exists $C(d)>0$ such that for all $u\geq 0$,
$\mathbb{P}\big(\big|m_{N}(z)-\mathbb{E}m_{N}(z)\big|\geq u\big)\leq C\exp\big(-N\eta^{4}u^{2}/C\big).$ (5.10)
Together with (5.9), this completes the proof. ∎
5.4. Conclusion
Proof of Lemma 3.4.
Let
$\widehat{\mu}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}}$
denote the empirical measure of $W_{N}$. By a basic property of Stieltjes transforms, the convergence of Stieltjes transforms in probability at every $z\in\mathbb{H}$ provided by Proposition 5.7 implies that $\widehat{\mu}_{N}$ converges weakly to $\nu$ in probability (see, e.g., [14, Lemma 2.1]). Since $\nu$ is absolutely continuous, by Lemma 5.5, we have that $\widehat{\mu}_{N}(I)$ converges to $\nu(I)$ in probability for every interval $I\subset\mathbb{R}$. The conclusion now follows from Lemma 5.5 and Lemma 5.6. ∎
@@ 6. Diagonal Resolvent Convergence
This section proves Lemma 3.5, which concerns convergence of the diagonal resolvent entries of $W_{N}$ to $R_{00}$. The two main technical ingredients of the proof are (i) a Combes–Thomas-type exponential decay bound for resolvents entries at fixed $\eta>0$ and (ii) the fact that a random $d$-regular graph is locally tree-like with high probability. Using these inputs, we prove Lemma 3.5 by truncating the $d$-regular graph around a fixed vertex and then controlling the differences of the resolvent entries of the original and truncated graph by the Combes–Thomas bound.
@@ 6.1. Preliminary Bounds
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be any (finite or infinite) graph with maximum degree at most $d$. Here $\mathcal{V}$ denotes the vertex set and $\mathcal{E}$ denotes the edge set. Let $A$ denote the adjacency operator on $\ell^{2}(\mathcal{V})$, let $V$ also denote a real diagonal potential (not necessarily bounded), fix $t>0$, and define H=tA+VH=-tA+V. Since this notation conflicts with the shorthand introduced in in Section 2.1 for the Anderson model on the Bethe lattice, we will always use the more precise notation $H_{d,t}=-tA^{(d)}+V^{(d)}$ when discussing the Bethe lattice in this section.
@@
Let $\mathcal{G}$ be a graph with maximum degree at most $d$. Fix $z=E+\mathrm{i}\eta$ with $\eta>0$ and set
@@ 6.2. Decoupling and Local Tree-Likeness
Let $K=H-H^{(r)}$. Then $K$ is supported on edges crossing the cut between $B_{r}(o)$ and $B_{r}(o)^{c}$, and it and equals t-t times the adjacency matrix of the graph formed by the cut edges. Hence, by Lemma 6.1, $\|K\|\leq t\|A\|\leq td$. It follows from the definitions of $G$ and $G^{(r)}$ that
$G-G^{(r)}=GKG^{(r)}.$
Let $\partial B_{r}(o)$ denote the set of vertices in $\mathcal{B}$ incident to at least one cut edge (both inside and outside $B_{r}(o)$). Then $\operatorname{dist}(\{o\},\partial B_{r}(o))=r$ and $K=\chi_{\partial B_{r}(o)}K\chi_{\partial B_{r}(o)}$. Therefore
@@
We then use local tree-likeness of random $d$-regular graphs to compare finite neighborhoods with the tree. For a rooted graph $(\mathcal{G},o)$ we write $B_{r}(\mathcal{G},o)$ for the rooted radius-$r$ ball.
@@
Let $G^{\mathrm{conf}}_{N,d}$ be the random $d$-regular multigraph from the configuration model (which places $d$ half-edges on of the $N$ vertices, then forms a perfect matching of the half-edges uniformly at random), and let $\mathbb{Q}$ be the induced measure on graph events. We let $C(d)>0$ denote a $d$-dependent constant that may increase at each appearance in this proof. For fixed $d$, we have $\mathbb{Q}(G^{\mathrm{conf}}_{N,d}\text{is simple})>C^{-1}$ for all $N\geq C$ by (4.5). Therefore, by the definition of conditional probability, we have for any event $\mathcal{E}$ measurable with respect to the graph and $N\geq C$ that
@@
We can construct $B_{r}(1)$ by breadth-first search in the associated configuration model, pairing one half-edge at each step. Let $|B_{r}(1)|$ denote the number of vertices in $B_{r}(1)$, and let $M$ be the total number of half-edges in $B_{r}(1)$ Deterministically,
@@
where we increase $C$ at each step if necessary. During breadth-first pairing procedure, at any step, the probability that a newly exposed half-edge pairs with one of the at most $M$ previously exposed half-edges is bounded by $CM/Nd$. A cycle in $B_{r}(1)$ can occur only if some such collision happens. By a union bound over the at most $M$ steps,

### 19 Mar 2026 v2

@@ 4.4. Conclusion
Fix $s\in(0,1)$ and $\delta\in(0,\eta^{-1}]$. There exists $C(d)>0$ such that for every $u>0$,
$\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\big]\big|\geq u\Big)\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Ct^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big).$ (4.11)
@@
Let $s\in(0,1)$ and $\delta\in(0,\eta^{-1}]$. Then for every $u>0$,
@@
Next, by Lemma 4.10 and Lemma 4.9, for any $s\in(0,1)$, $\delta\in(0,\eta^{-1}]$, and $u>0$,
@@
$\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Ct^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big),$
@@ 5. Diagonal Resolvent Convergence
This section proves Lemma 3.4, which concerns convergence of the diagonal resolvent entries of $W_{N}$ to $R_{00}$. The two main technical ingredients of the proof are (i) a Combes–Thomas-type exponential decay bound for resolvent entries at fixed $\eta>0$ and (ii) the fact that a random $d$-regular graph is locally tree-like with high probability. Using these inputs, we prove Lemma 3.4 by truncating the $d$-regular graph around a fixed vertex and then controlling the differences of the resolvent entries of the original and truncated graph by the Combes–Thomas bound.
@@ 5.1. Preliminary Bounds
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be any (finite or infinite) graph with maximum degree at most $d$. Here $\mathcal{V}$ denotes the vertex set and $\mathcal{E}$ denotes the edge set. Let $A$ denote the adjacency operator on $\ell^{2}(\mathcal{V})$, let $V$ also denote a real diagonal potential (not necessarily bounded), fix $t>0$, and define H=tA+VH=-tA+V. Since this notation conflicts with the shorthand introduced in Section 2.1 for the Anderson model on the Bethe lattice, we will always use the more precise notation $H_{d,t}=-tA^{(d)}+V^{(d)}$ when discussing the Bethe lattice in this section.
@@
Let $\mathcal{G}$ be a graph with maximum degree at most $d$. Fix $z=E+\mathrm{i}\eta\in\mathbb{H}$ and set
@@ 5.2. Decoupling and Local Tree-Likeness
Let $K=H-H^{(r)}$. Then $K$ is supported on edges crossing the cut between $B_{r}(o)$ and $B_{r}(o)^{c}$, and it equals t-t times the adjacency matrix of the graph formed by the cut edges. Hence, by Lemma 5.1, $\|K\|\leq t\|A\|\leq td$. It follows from the definitions of $G$ and $G^{(r)}$ that
$G-G^{(r)}=-GKG^{(r)}.$
Let $\partial B_{r}(o)$ denote the set of vertices in $\mathcal{V}$ incident to at least one cut edge (both inside and outside $B_{r}(o)$). Then $\operatorname{dist}(\{o\},\partial B_{r}(o))=r$ and $K=\chi_{\partial B_{r}(o)}K\chi_{\partial B_{r}(o)}$. Therefore
@@
We now use local tree-likeness of random $d$-regular graphs to compare finite neighborhoods with the tree. For a rooted graph $(\mathcal{G},o)$ we write $B_{r}(\mathcal{G},o)$ for the rooted radius-$r$ ball.
@@
Let $G^{\mathrm{conf}}_{N,d}$ be the random $d$-regular multigraph from the configuration model (which places $d$ half-edges on each of the $N$ vertices, then forms a perfect matching of the half-edges uniformly at random), and let $\mathbb{Q}$ be the induced measure on graph events. We let $C(d)>0$ denote a $d$-dependent constant that may increase at each appearance in this proof. For fixed $d$, we have $\mathbb{Q}(G^{\mathrm{conf}}_{N,d}\text{is simple})>C^{-1}$ for all $N\geq C$ by (4.5). Therefore, by the definition of conditional probability, we have for any event $\mathcal{E}$ measurable with respect to the graph and $N\geq C$ that
@@
We can construct $B_{r}(1)$ by breadth-first search in the associated configuration model, pairing one half-edge at each step. Let $|B_{r}(1)|$ denote the number of vertices in $B_{r}(1)$, and let $M$ be the total number of half-edges in $B_{r}(1)$. Deterministically,
@@
where we increase $C$ at each step if necessary. During the breadth-first pairing procedure, at any step, the probability that a newly exposed half-edge pairs with one of the at most $M$ half-edges attached to already discovered vertices is bounded by $CM/Nd$. A cycle in $B_{r}(1)$ can occur only if some such collision happens. By a union bound over the at most $M$ steps,